Minutes of the 20th SOHO SWT Meeting

Goddard Space Flight Center, Greenbelt, MD 15 August 1996 Contents

1 Action Items SOHO SWT–20 2 2 Agree Agenda and Actions Revision 2 3 S/C Status 2 4 Telemetry Redistribution 2 5 New Significant Results 3 6 Science Priorities 3 7 Community involvement in SOHO 3

7.1 Guest Investigators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 7.2 Operations participation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 7.3 Publication policies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 7.4 Mission after March 1998 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

8 Archives 4 9 PR matters 4 10 SOHO Workshops and Colloquia 4 11 AOB 4 Annex 1: Agenda 5 Annex 2: List of Participants 6 Annex 3: Spacecraft Status 7 Annex 4: EIT Telemetry 18 Annex 5: Status of Experiments 23 Annex 6: GI Programme 36 Annex 7: Data Issues 40

1 Action Items SOHO SWT–20

Action 20-1: On PIs to determine if the new computation of the spacecraft velocity in heliocentric coordinates is correct. (Samples can be found at http://sohowww.nascom.nasa.gov/ data/ancillary/orbit/definitive/new samples). Input to L. Sanchez by September 6.

Action 20-2: On ESA to provide a detailed script for the station keeping/momentum management maneuver on 11-12 September, including window for SSU/SEU patch upload and testing. Due date: 30 August.

Action 20-3: On ESA/Matra, MDI and VIRGO to analyse S/C pointing disturbances.

Action 20-4: On PS to establish Working Group to study technical matters of telemetry redistribution with relevant parties (EIT, CDS, SUMER, UVCS, S/C).

Action 20-5: On PS to organize series of small workshops at SOHO EAF.

2 Agree Agenda and Actions Revision

Agenda — annex 1 List of Participants — annex 2

Actions revision

Action 19-1: ongoing On MDI, SUMER, EIT/LASCO, UVCS to provide each month solar center coordinates for assessment of medium term pointing stability. Input to PS by 20th of each month.

Action 19-2: closed On PS to request reprocessing effort of all level 0 data from tapes. Due immediately.

Action 19-3: closed On PIs to revisit number of CD-ROM copies they need. Input to PS before 15 June.

3 S/C Status

Presentation by J.-P. Olive — see annex 3. Next orbit correction maneuver and momentum management planned for 12-13 September. SSU/SEU patch on September 11. More details and exact schedule to be provided by ESA before 30 August (Action 20-2). There is an action on ESA/Matra, MDI and VIRGO to analyse S/C pointing disturbances in order to understand the source of the elusive 8 hour periodicities found in VIRGO/LOI data (Action 20-3).

4 Telemetry Redistribution

J. Gurman presented what EIT can do now with its telemetry allocation of 1 kb/s, and what they would like to be able to do (do more science, especially joint science with other SOHO instruments – see annex 4). He then presented a “modest proposal”, which in essence would triple the EIT bandwith to 3 kb/s. In practical terms, this means donations totalling 2.0 kb/s of telemetry. EIT’s position was well appreciated by all attendees. There is an action on the PS to establish a Working Group to study in detail the technical matters of telemetry redistribution with the relevant parties (EIT, CDS, SUMER, UVCS, S/C) (Action 20-4).

5 New Significant Results

See forthcoming papers in Solar Physics, COSPAR, ... For MDI 2 months data coverage see annex 5.

6 Science Priorities

i) CMEs ii) Flows in coronal loops iii) Large scale structure of the solar minimum corona (repeat JOP044: Whole Sun Month

in December; tentative starting date: 9 December) iv) Low latitude coronal holes (target of opportunity)

7 Community involvement in SOHO

7.1 Guest Investigators

— see annex 6

• About 100 Letters of Intent received.

• Proposals are due on 20 September.

• The PIs will advise on the proposals that require data from their instrument.

• Proposals will be declared acceptable or not by the GI selection committee (GISC) on the basis of their scientific merit and feasibility; the GISC will meet on 20-22 November 1996.

• NASA will decide on the funding of US proposals according to available funds.

• Non-US proposers are responsible for securing their own funding.

PIs are requested to comment on “Proposal advise by PI” form (see annex 6).

7.2 Operations participation

Art Poland expressed his concerns about the participation in science operations and analysis at the EOF and EAF. A lively discussion started. One of the suggestions made to bring more people in was to organize a series of small, dedicated workshops. Topics: Polar plumes, the Whole Sun Month campaign, XBPs, etc.. Suggestions are most welcome! Project Scientists to organize (Action 20-5).

7.3 Publication policies

SUMER and CDS have submitted their first results papers for publication in Solar Physics. For them the “flood gates are open” and the community is free to publish papers using CDS and SUMER data, subject to the data rights policy. Apparently the situation is different for the other teams. V. Domingo conveyed Martin Huber’s appreciation for the PI’s help by making material available for inclusion into his COSPAR Opening Lecture. He has now been asked to write up this lecture for the COSPAR Information Bulletin (CIB), for which he would like to have the PIs concurrence. No objections were brought forward. Instead, a repeated exposure of SOHO to the COSPAR members through the CIB was considered a good PR effort and beneficial for the mission’s recognition in the space community at large.

7.4 Mission after March 1998

Has not been discussed (premature).

8 Archives

L. Sanchez presented the status of the SOHO archive and catalogues (see annex 7). The first SOHO summary data CD will be produced for the month of August, with previous months published later. First month, for which a summary data CD will be produced will be May 1996 (start of nominal mission). Campaign catalogue: input needs to be made before run at weekly meeting to get a campaign number assigned to JOP. SOCs to ensure proper entries. There is an action on the PIs to determine if the new computation of the spacecraft velocity in heliocentric coordinates is correct. (Samples can be found at http://sohowww.nascom.nasa. gov/data/ancillary/orbit/definitive/new samples). Input to L. Sanchez by September 6.

9 PR matters

Next (tentative) event: Space Science Update on results from MDI/SOI (coordinated with a release in Europe) targeted for 15 October 1996 (after Nice meeting).

In order to promote visibility for SOHO and possibly make it a regular feature of earth weather forecasts, Fabricio Felici suggested the possibility of issuing to meteo offices and/or major broadcast corporations monthly or bi-monthly an EIT and/or LASCO movie of the preceding period with a brief description of the evolution of the Sun’s behaviour, a sort of “weather summary” of the Sun. This idea was very well received by the SWT and it was decided to go ahead with it. The plan is to provide a first proposal some time in September, after consultation of PR specialists.

10 SOHO Workshops and Colloquia

• SOHO V: “The quiet Sun atmosphere and wind”: Oslo, 10-13 June 1997

• 33rd ESLAB Symposium “Solar Events and their Manifestation in the Heliosphere and in Geospace” (Results from the 3rd and 4th IACG Campaigns): ESTEC, 23-25 September 1997

• SOHO VI “Helioseismology” (with GONG ’97): Fall 1997, Harvard(?)

11 AOB

Next SWT meeting: 14/15 November 1996 at GSFC (again in conjunction with the monthly SPWG meeting); Science Workshop postponed to Friday afternoon and Saturday (due to Ulysses SWT earlier this week at JPL).

In order to help better integrate the remote PI team members in the SOHO community it was proposed to have future SWT meetings also at PI home institutes. PIs interested in hosting the February ’97 meeting should contact Vicente Domingo.

Next S/C roll for SWAN sensor crosscalibration and JOP002 in conjunction with the next momentum management (presumably in November). Duration: ≤ 16 hours. If possible the roll should be done prior to the momentum management, not after.

Annex 1: Agenda

5

Annex 2: List of Participants

6

Annex 3: Spacecraft Status

7

Annex 4: EIT Telemetry

18

Annex 5: Status of Experiments

23

Annex 6: GI Programme

36

Annex 7: Data Issues

40